OpenStudy (anonymous):

The perimeter of a rectangle is 124 inches. The length exceeds the width by 10 inches. what is the length in inches? what is the width in inches?

OpenStudy (anonymous):

26 and 36

OpenStudy (saifoo.khan):

let widht be x let lenght b (x+10) finally 2(l+ b) = 124

OpenStudy (tad1):

perimeter of a rectangle is 2x + 2y = 124 y = w + 10 substitute 2x + 2 (x + 10) = 124 solve for x. Use x to back substitute for y 2x + 2x + 20 = 124 4x = 104 x = 26 y = 36 check 26*36 = 124 i hope

OpenStudy (saifoo.khan):

2 [x+10(x)] 2 [x^2 + 10x] = 124

OpenStudy (anonymous):

a+b = 62 and a-b = 10 => a=36, b = 26

OpenStudy (anonymous):

so length is 36 and width is 26?

OpenStudy (saifoo.khan):

yes!

OpenStudy (anonymous):

thank you all